# Portability of Fortran’s ‘do concurrent’ on GPUs

### v1

@@ II Related Work
The most computationally expensive kernels of the quantum chemistry production code GAMESS were tested using DC alone on CPUs and GPUs in [9]. The authors discovered that running the DC implementation on a single GPU surpassed the performance of their earlier OpenACC and OpenMP target implementations.
@@
Recently, DC was used to GPU-accelerate the production relativistic MHD astrophysics code ‘ECHO’ [11]. The authors relied on automatic managed memory, but still had to use some OpenACC directives for the code to work correctly (mostly dealing with subroutines within DC loops). They demonstrated impressive scaling up to 256 quad-GPU nodes in cases where the computation time constituted a much larger fraction of the run time compared to MPI communication.
@@ III High Performance Flux Transport (HipFT)
[…]
@@ III-A Parallel implementation
An often difficult code algorithm to parallelize on GPUs are array reductions [18]. An example array reduction in HipFT is shown in Listing 2.
@@
fn2_fn1 = zero
fs2_fs1 = zero
do concurrent(k=2:npm-1) reduce(+:fn2_fn1,fs2_fs1)
fn2_fn1 = fn2_fn1 &
+ (diffusion_coef(1 ,k,i) &
@@
Loops like these can be difficult for the compiler to efficiently parallelize (see Sec. VI).
@@ IV-A NVIDIA
NVIDIA has the most mature implementation of DC for GPU offload. The NV HPC SDK compiler nvfortran (NV) version 24.5 is used to compile the code. Under the hood, the compiler uses its OpenACC back-end to compile the DC loops. As discussed above, this requires us to add an OpenACC device selection directive in addition to the equivalent OpenMP directive.
@@ IV-B INTEL
Support for DC GPU offload is a recent addition (introduced in version 2024.0) to the Intel compiler ifx and is implemented using an OpenMP target back-end. The compiler does not currently have an automatic memory management system for DC, so to avoid the significant performance reduction from GPU-CPU transfers, we use manual memory management using pairs of OpenMP target enter/exit data directives. These pairs form unstructured data regions, which allows multiple data entry and exit points across different regions of the code. In the original Intel implementation, the DC loops were transformed into OpenMP target loop directives with an ‘always’ map-type-modifier indicating to always transfer the data to and from the GPU. In the latest release (as of this writing, 2024.2), a special maptype compiler flag (shown in Listing IV-B) has been added to let the programmer specify the map-type-modifier for all transformed OpenMP target regions. Setting this value to ”present” tells the compiler to assume all data (other than scalars) in DC loops are already on the device, leaving the memory management for the user to take care of using OpenMP target data directives.
@@
In our preliminary GPU tests, we found that a few specific loops were running far more slowly than expected. These were the array reductions with nested DC loops as shown in Sec. III-A. After investigation, the compiler team found that adding !\$omp parallel loop above the inner DC loops changes the compiler’s parallelization and greatly increases performance (see Sec.VI). While adding these directives is not standard in a DC loop (and adding them can cause a compiler error using other compilers) they work here due to the Intel compiler using OpenMP as the DC back-end. We note that the compiler team is working on updates that should make adding these directives unnecessary. We indicate this small (6 line) modification of the code in the results with an asterisk (*).
@@ IV-C AMD
where -homp and -hacc activates the OpenMP and OpenACC runtimes respectively, -hcpu and -haccel set the CPU and GPU hardware targets respectively, and -hoffload_do_concurrent activates the DC GPU offload. It is also necessary to set the environment variable HSA_XNACK=1 to activate the automatic unified memory management [^11] https://rocm.docs.amd.com/en/latest/conceptual/gpu-memory.html.
@@ V Test case
For each compiler/GPU/CPU tested, we first ensure correct code execution by making sure the testsuite included with the HipFT code passes. For the main results, we use the included example run flux_transport_1rot_flowAa_diff_r8 as our test case. This run starts with a pole-filled, processed full-Sun ‘synoptic’ map for Carrington rotation 2252 at a resolution of 1024x512. It simulates surface flux transport (using both advection and diffusion) for 28 days (approximately one solar rotation) using analytically prescribed meridional and differential rotation flows. Eight realizations are computed (allowing up to 8 MPI ranks) where four levels of diffusion are used, and a magnetic field based flow attenuation is or is not applied, leading to the eight total realizations. A visualization of the run is shown in Fig. 1.
@@
As mentioned in Sec. IV, HipFT uses OpenMP target directives for manually managing the GPU memory. For NVIDIA, using these directives is optional, so we run the test case with manual memory, managed memory, and unified memory where available. We note that when using managed or unified memory, the code becomes pure standard Fortran (with the exception of a single device selection directive). For Intel GPUs, we must use the manual data directives for efficient computation, while for AMD GPUs with CCE, we are required to rely on automatic memory management (standard Fortran).
@@ VI-A Data center / Server
As mentioned in Sec. IV-B, the current (as of this writing 2024.2) Intel compiler does not efficiently parallelize some nested DC loops. The problem has been identified and planned to be remedied. In the meantime, by adding !\$omp parallel loop to the inner nested DC loops, the current compiler efficiently parallelizes the loops. We denote this small (6-line) modification to the code with an asterisk (*). In Fig. 2, we show how this modification improves the performance on MAX 1100 and MAX 1550 GPUs.
Figure 2: Run time comparison (less is better) of the HipFT test case on Intel data center GPUs between the original HipFT code and the modified code (adding !\$omp parallel loop to the inner nested DC loops). We see a substantial performance improvement with the modified code.
@@
Although we had to (temporarily) slightly modify the code to get good performance on the Intel GPUs and used OpenMP target for data movement, we were still able to use standard Fortran for the vast majority of the code. The performance of the Intel MAX 1550 GPU is quite good considering previous difficulties in achieving maximum performance [19]. This may be due to running HipFT with 2 MPI ranks, one per GPU tile. Using more MPI ranks typically improves performance on HipFT by spreading the problem out more evenly (for example, to yield best performance on the CPUs, we use the maximum number of MPI ranks possible spread evenly across NUMA domains). We note that we did not test using multiple MPI ranks within a single NVIDIA GPU using MPS [^12] https://docs.nvidia.com/deploy/mps or MIG [^13] https://www.nvidia.com/en-us/technologies/multi-instance-gpu. This is left for future testing.
@@ VI-B Consumer
We see that the upwind advection is many times faster than the WENO3 scheme due to the absence of divisions (on other GPUs, the upwind scheme is typically $1010$ to $3030$ times faster, not the $120120$ times we see here). As future Intel consumer GPUs may have hardware double precision compute units, in order to have a more representative comparison, we run the test case with the upwind method exclusively across all consumer platforms. Note that this makes these consumer results not comparable to the server results in Sec. VI-A.
@@
We see that the Intel Arc GPU does not perform at the expected level based on its memory bandwidth as expected, most likely due the double precision emulation. However, it is still capable of reasonable run times for the test case (given the upwind algorithm), outperforming many modern consumer CPUs.
@@ VII Summary and outlook
Three concepts that the Fortran language is lacking that are needed/wanted for GPU computing are the recognition and management of separate memory spaces, the recognition and selection of multiple devices, and asynchronous loop execution. With the advent of unified GPU-CPU architectures like the NVIDIA Grace-Hopper and AMD MI300A, along with automatic memory management capabilities of compilers, the data management is taken care of, and with alternate methods of launching the code, the device selection is as well. For discrete GPUs, manually managing the GPU-CPU data movement can be important for performance. Therefore, augmenting standard Fortran with data directives (either OpenMP target or OpenACC) is often preferred, and for the Intel platform is currently required for good performance.
In this paper, we were able to achieve good performance on NVIDIA and Intel GPUs (with some slight modifications) and showed that the performance difference between automatic memory management and manual data movement using OpenMP target directives on platforms that support both was minimal. We found that using automatic memory management on discrete GPUs slightly reduced performance, but when using it on a unified CPU-GPU SoC (in this case the GH200), the performance actually slightly increased. We were also able to compile and correctly run the code on AMD GPUs, but more development is needed for performance.
We also tested the code on consumer GPUs where possible, and showed good performance, especially for NVIDIA GPUs. Reasonable performance was also achieved for some algorithms on an Intel Arc GPU even though they are not designed for HPC workloads and require the use of double precision emulation. We expect that the performance on Intel’s upcoming consumer GPUs will be much better as they may have hardware double precision units. We were not able to test performance on consumer AMD GPUs, but we expect support for those is forthcoming.
[…]
@@ VIII Roofline plots of tested hardware
Here we use the OpenCL-Benchmark [21] to produce practical (instead of peak theoretical) roofline plots of the GPUs tested in this paper. We use the benchmark’s floating point and memory-bound results to show the best FLOP/s expected given a kernel/algorithm’s arithmetic intensity (FLOP/s per byte). The range of arithmetic intensities used in HipFT is highlighted. The server/data center results are shown in Fig. 6, while those for consumer hardware are shown in Fig. 7.

### v2

@@ II Related Work
The most computationally expensive kernels of the quantum chemistry production code GAMESS were tested using DC alone on CPUs and GPUs in [9]. The authors discovered that running the DC implementation on a single GPU surpassed the performance of their earlier OpenACC and OpenMP target implementations. This work was recently extended to the LibERI library interfaced to GAMESS, where the performance of the DC code outperformed their OpenMP Target implementation [liberi_gamess_dc].
@@
DC was also used to GPU-accelerate the production relativistic MHD astrophysics code ‘ECHO’ [11]. The authors relied on automatic managed memory, but still had to use some OpenACC directives for the code to work correctly (mostly dealing with subroutines within DC loops). They demonstrated impressive scaling up to 256 quad-GPU nodes in cases where the computation time constituted a much larger fraction of the run time compared to MPI communication.
@@ III High Performance Flux Transport (HipFT)
HipFT implements advection, diffusion, and data assimilation and can compute multiple realizations in a single run across model parameters to create ensembles of full-Sun maps. It employs finite difference methods on a spherical grid and utilizes highly accurate and high-performance numerical methods, such as a Strong Stability Preserving Runge-Kutta (SSPRK(4,3) [14]) scheme with a Weighted Essentially Oscillatory (WENO3-CS(h) [15]) scheme for advection, and the Extended Stability Runge-Kutta Gegenbauer Super Time Stepping (RKG2(3/2) [16]) scheme for diffusion. Computationally, these methods boil down to memory-bound array and stencil operations. HipFT is written in modern Fortran ($\sim 9000$ lines of code) and parallelized for use with multiple multi-core CPU and multi-GPU compute nodes. Details of the code are described in Ref. [17].
@@ III-A Parallel implementation
An often difficult algorithm to parallelize on GPUs are array reductions (see [18]). Depending on the order of the loops, either an atomic operation is needed for each reduction array element (which may not be efficient), or several scalar reductions are required, which may be serialized unless the topology of the GPU parallelism is optimally chosen by the compiler (such as spreading the array across compute blocks, and using threaded parallelism within each block with shared memory). In HipFT, the array reductions use the second approach. An example of an array reduction within the code is shown in Listing 2.
@@
fn = zero
fs = zero
do concurrent(k=2:npm-1) reduce(+:fn,fs)
fn = fn + (diffusion_coef(1 ,k,i) &
@@ IV-A NVIDIA
NVIDIA has the most mature implementation of DC for GPU offload. The NV HPC SDK compiler nvfortran (NV) version 24.5 is used to compile the code. Under the hood, the compiler uses its OpenACC backend to compile the DC loops. As discussed above, this requires us to add an OpenACC device selection directive in addition to the equivalent OpenMP directive.
@@ IV-B INTEL
Support for DC GPU offload is a recent addition (introduced in version 2022.0) to the Intel compiler ifx and is implemented using an OpenMP target backend. The compiler does not currently have an automatic memory management system for DC, so to avoid the significant performance reduction from GPU-CPU transfers, we manually manage the memory using pairs of OpenMP target enter/exit data directives. These pairs form unstructured data regions, which allows multiple data entry and exit points across different regions of the code. In the original Intel implementation, the DC loops were transformed into OpenMP target loop directives with an ‘always’ map-type-modifier indicating to always transfer the data to and from the GPU. In the latest release (as of this writing, 2024.2), a special maptype compiler flag (shown in Listing IV-B) has been added to let the programmer specify the map-type-modifier for all transformed OpenMP target regions. Setting this value to ”present” tells the compiler to assume all data (other than scalars) in DC loops are already on the device, leaving the memory management for the user to take care of using OpenMP target data directives.
@@
In our preliminary GPU tests, we found that a few specific loops were running far slower than expected. These were the array reductions with nested DC loops as shown in Sec. III-A. After investigation, it was found that adding !\$omp parallel loop above the inner DC loops changes the internal heuristics of the compiler, resulting in a far more efficient kernel (see Sec.VI). Adding these directives is not standard in a DC loop (and can cause an error when using other compilers). However, the modified heuristics are being integrated into the default compilation, so future compiler releases will not require these added directives. We indicate this small (6 lines) modification of the code in the results with an asterisk (*).
@@ IV-C AMD
where -homp and -hacc activates the OpenMP and OpenACC runtimes respectively (both are necessary as CCE uses an OpenACC backend for DC), -hcpu and -haccel set the CPU and GPU hardware targets respectively, and -hoffload_do_concurrent activates the DC GPU offload. It is also necessary to set the environment variable HSA_XNACK=1 to activate the automatic unified memory management [^11] https://rocm.docs.amd.com/en/latest/conceptual/gpu-memory.html.
@@ V Test case
For each compiler/GPU/CPU tested, we first ensure correct code execution by making sure the testsuite included with the HipFT code passes. The testsuite validation criteria uses various derived quantities of the solution which are checked to be within a relative tolerance of $1\times 10^{-5}$ compared to a reference solution.
For the main results presented here, we use the included example run flux_transport_1rot_flowAa_diff_r8 as our test case. This run starts with a pole-filled, processed full-Sun ‘synoptic’ map for Carrington rotation 2252 at a resolution of 1024x512. It simulates surface flux transport (using both advection and diffusion) for 28 days (approximately one solar rotation) using analytically prescribed meridional and differential rotation flows. Eight realizations are computed (allowing up to 8 MPI ranks) where four levels of diffusion are used, and a magnetic field based flow attenuation is or is not applied, leading to the eight total realizations. A visualization of the run is shown in Fig. 1.
@@
As mentioned in Sec. IV, HipFT uses OpenMP target directives for manually managing the GPU memory. For NVIDIA, using these directives is optional, so we run the test case with manual memory, managed memory, and unified memory where available. We note that when using managed or unified memory, the code becomes pure standard Fortran (with the exception of a single device selection directive). For Intel GPUs, we must use the manual data directives for efficient computation (as unified memory is not yet supported), while for AMD GPUs with CCE, we are required to rely on automatic memory management (as directive-based data management is not currently supported by CCE).
@@ VI Results
As HipFT was developed as a DC-based code from the ground up, we do not compare the performance of individual DC loops to alternative GPU-offload methods (such as OpenACC/MP, CUDA-Fortran, etc.). For examples of such comparisons, see Ref. [5] and [10], where the DC performance was found to be similar to other forms of GPU-offload. Our focus here is the portability of DC GPU offload.
@@ VI-A Data center / Server
As mentioned in Sec. IV-B, the current (as of this writing 2024.2) Intel compiler does not efficiently parallelize some nested DC loops. The problem has been identified and planned to be remedied. In the meantime, by adding !\$omp parallel loop to the inner nested DC loops, the current compiler efficiently parallelizes the loops. We denote this small (6-line) modification to the code with an asterisk (*). In Fig. 2, we show how this modification improves the performance on MAX 1100 and MAX 1550 GPUs. The MAX 1550 GPU is designed with two tiles which can be set up to appear as a single GPU (known as ”composite” mode) or as two separate GPUs (”flat” mode) [^12] https://spec.oneapi.io/level-zero/latest/core/PROG.html#device-hierarchy. For all results in this paper, we use the full GPU with the 2 tiles seen as separate GPUs (1 MPI rank each) as we found the performance is better than using them together as one device.
Figure 2: Run time comparison (less is better) of the HipFT test case on Intel data center GPUs between the original HipFT code and the modified code (adding !\$omp parallel loop to the inner nested DC loops). We see a substantial performance improvement with the modified code (for these results, the file I/O time has been omitted from the ”Other” category).
@@
Although we had to (temporarily) slightly modify the code to get good performance on the Intel GPUs and use OpenMP target for data movement, we were still able to use standard Fortran for the vast majority of the code. The performance of the Intel MAX 1550 GPU is quite good considering previous difficulties in achieving maximum performance [19]. This may be due to running HipFT with 2 MPI ranks, one per GPU tile. Using more MPI ranks typically improves performance on HipFT by spreading the problem out more evenly (for example, to yield best performance on the CPUs, we use the maximum number of MPI ranks possible spread evenly across NUMA domains). We note that we did not test using multiple MPI ranks within a single NVIDIA GPU using MPS [^13] https://docs.nvidia.com/deploy/mps or MIG [^14] https://www.nvidia.com/en-us/technologies/multi-instance-gpu. This is left for future testing.
@@ VI-B Consumer
We see that the upwind advection is many times faster than the WENO3 scheme due to the absence of divisions (on other GPUs, the upwind scheme is typically $10$ to $30$ times faster, not the $120$ times we see here). As future Intel consumer GPUs may have hardware double precision compute units, in order to have a more representative comparison, we run the test case with the upwind method exclusively across all consumer platforms. Note that this makes these consumer results not comparable to the server results in Sec. VI-A.
@@
We see that the Intel Arc GPU does not perform at the expected level based on its memory bandwidth as expected, most likely due to the double precision emulation. However, it is still capable of reasonable run times for the test case (given the upwind algorithm), outperforming many modern consumer CPUs.
@@ VII Summary and outlook
Three concepts that the Fortran language is lacking that are needed/wanted for GPU computing are the recognition and management of separate memory spaces, the recognition and selection of multiple devices, and asynchronous loop execution. With the advent of unified GPU-CPU architectures like the NVIDIA Grace-Hopper and AMD MI300A, along with automatic memory management capabilities of compilers, the data management is taken care of. With alternate methods of launching the code (see Ref. [10] for an example), the device selection can be done without directives as well. For discrete GPUs, manually managing the GPU-CPU data movement can be important for performance. Therefore, augmenting standard Fortran with data directives (either OpenMP target or OpenACC) is often preferred, and for the Intel platform is currently required for good performance.
In this paper, we were able to achieve good performance on NVIDIA and Intel GPUs (with some slight modifications) and showed that the performance difference between automatic memory management and manual data movement using OpenMP target directives on platforms that support both was minimal. We found that using automatic memory management on discrete GPUs slightly reduced performance, but when using it on a unified CPU-GPU SoC (in this case the GH200), the performance actually slightly increased. We were also able to compile and correctly run the code on AMD GPUs, but more development is needed for good performance.
We also tested the code on consumer GPUs where possible, and showed good performance, especially for NVIDIA GPUs. Reasonable performance was also achieved for some algorithms on an Intel Arc GPU even though it is not designed for HPC workloads and requires the use of double precision emulation. We expect that the performance on Intel’s upcoming consumer GPUs will be much better as they may have hardware double precision units. We were not able to test performance on consumer AMD GPUs, but we expect support for those is forthcoming.
We note that HipFT is not a large code ($\sim 9000$ lines) and does not use certain features that are known to be challenging for GPU acceleration (such as GPU-aware MPI, derived type arrays, and function calls within parallel regions [18]). We plan to continue testing DC across vendors for larger codes that do contain those features.
@@ VIII Roofline plots of tested hardware
Here we use the OpenCL-Benchmark [21] to produce roofline plots of the GPUs tested in this paper. The optimized and simple computations of the benchmark make it ideal as a practical (instead of peak theoretical) way to generate roofline plots for comparison. We use the benchmark’s floating point and memory read/write results to show the best FLOP/s expected given a kernel/algorithm’s arithmetic intensity (FLOP/s per byte). The floating point results in the OpenCL-Benchmark are computed with fused multiply adds and appear as
kernel_double(global float* data){
double x = (double)get_global_id(0);
double y = (double)get_local_id(0);
for(uint i=0u;i<128u;i++){
x = fma(y, x, y);
y = fma(x, y, x);}
data[get_global_id(0)] = (float)y;}
while the coalesced read and write kernels are computed as
kernel_coalesced_write(global float* data){
const uint n = get_global_id(0);
for(uint i=0u;i<def_M;i++) data[i*def_N+n]=0.0f;}
kernel_coalesced_read(global float* data){
const uint n = get_global_id(0);
float x = 0.0f;
for(uint i=0u;i<def_M;i++) x += data[i*def_N+n];
data[n] = x;}
The server/data center results are shown in Fig. 6, while those for consumer hardware are shown in Fig. 7. In each, the range of arithmetic intensities used in HipFT is highlighted.
@@
IX Artifact Availability Statement
Artifact Identification
We have investigated the portability of using Fortran’s ‘do concurrent‘ (DC) for accelerated computing by testing its use across GPU vendors using the solar flux evolution code HipFT. HipFT is written using DC loops and includes OpenMP target data directives for manual data movement between the CPU and GPU for platforms that do not support (or have non-optimal) automatic data management. We successfully compiled and ran a production-level test case on NVIDIA, Intel, and AMD GPUs using NVIDIA’s HPC SDK’s nvfortran, Intel’s OneAPI HPC Toolkit’s ifx, and HPE’s Cray Compiler Environment’s ftn compilers respectively. Both server and consumer hardware were tested. Tags/releases of the github repository of HipFT (https://github.com/predsci/hipft) were made for the versions used in this paper, specifically the v1.11.0 release of the code was used for CPU runs and the NVIDIA GPU and AMD GPU runs. The v1.11.0-intel-gpu branch release (only 6 modified lines) was used for Intel GPU runs. The test case is the run located in the examples/flux_transport_1rot_flowAa_diff_r8 folder of the HipFT github repository. Given access to the computational hardware, the repository with its release tags can be used to directly reproduce the run timings shown in the paper by building with the compiler flags described in the paper, along with any/all system-specific compiler flags that may be needed. The HipFT code outputs a breakdown of the time taken in each section of the code, and this data is used to generate the bar plots in the figures.
Reproducibility of Experiments
The workflow to generate the results in this paper is as follows: First, checkout the relevant release of HipFT (either github.com/predsci/HipFT/releases/tag/v1.11.0 for NVIDIA and AMD GPUs or github.com/predsci/HipFT/releases/tag/v1.11.0-intel-gpu for Intel GPUs). Next, build the code according to the build instructions (including having an HDF5 library installed with the same compiler used for HipFT), the included sample build scripts, and the compiler flags described in the paper. Then, copy the input files from the examples/flux_transport_1rot_flowAa_diff_r8 folder into a new folder, and run the code with mpirun -np <N> hipft (assuming HipFT is in your run path) and where <N> is the number of MPI ranks to use (in this case, typically 1).
The execution time will depend on the hardware and software platforms. If running on similar hardware as described in the paper, the run times should be very close to those in the bar plots (maximum $\sim 1$hr), except possibly for the gray ”other” area which can be sensitive to the speed of the file system and network.
If the test case is run on the same hardware and software versions (including compilers) as done in the paper, it is expected to yield very similar run times as those shown in the bar plots. The results should show the portability of Fortran’s DC for GPU acceleration, with similar timings for NVIDIA and Intel GPUs of comparable capabilities, as well as correct execution on AMD GPUs. Running the tests with newer versions of the compilers, or on newer GPU hardware should result in faster run times.